Can you help me with this problem (a+1)^2=25
Found 2 solutions by harpazo, Alan3354:
Answer by harpazo(655)

Apply FOIL on the left side.
(a+1)(a+1) = 25
a^2+2a+1 = 25
a^2+2a+1 - 25 = 0
a^2+2a-24 = 0
(a-4)(a+6) = 0

Set each factor to equal zero and solve for a.
a-4 = 0
a = 4

Now the next factor.
a+6 = 0
a = -6

Answer: a = 4 and a = -6.



Answer by Alan3354(69443)

Can you help me with this problem (a+1)^2=25
-----
Get the square root of each side.
a+1 = 5
a = 4
============
a+1 = -5
a = -6